OpenStudy (anonymous):

4.16^2x=124 2x*ln(16) = ln(124)

OpenStudy (anonymous):

4.16^2x=124 1.0.869 2.1.150 3.1.739 4.3.477

OpenStudy (anonymous):

lol didn't I write the 2x*ln(16) = ln(124)

OpenStudy (anonymous):

nope. i copied what you wrote.

OpenStudy (anonymous):

just solve for x

OpenStudy (anonymous):

54

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

what did i do wrong

OpenStudy (anonymous):

no idea, show me the steps and I will help you

OpenStudy (anonymous):

16-124=108 /2 =-54

OpenStudy (anonymous):

\[2x*\ln(16) = \ln(124)\] \[2x=\frac{ \ln(124) }{ \ln(16) } \] \[x=\frac{ \ln(124) }{ 2\ln(16) } \]